A. 0.191 g of steel was dissolved in acid and diluted to 100 mL with water. The solution was found to contain 7.7 ppm Mn by atomic absorption analysis. What was the weight percent Mn in the steel?

B. A multivitamin tablet with a label value of 13 mg Zn per tablet was dissolved in acid and diluted to 100 mL with water. What is the theoretical Zn concentration of the solution? Answer in ppm.
Answer #1

A

We have ppm = Mass of solute / Mass of solution x 10 6

Assume density of solution = 1 g / ml then mass of solution = 100 g

Therefore, Mass of Mn / 100 g x 10 6 = 7.7

Mass of Mn / 100 g = 7.7 / 10 6 = 7.7 x 10 -06

Mass of Mn = 7.7 x 10 -06 x 100 g = 7.7 x 10 -04 g

i e 100 ml diluted solution contain 7.7 x 10 -04 g Mn

Therefore, 0.191 g of steel contain 7.7 x 10 -04 g Mn

% of Mn in steel = 7.7 x 10 -04 g / 0.191 x 100

= 0.403

B ) We have ppm = Mass of solute / Mass of solution x 10 6

Mass of Zn = 12 mg = 0.012 g

ppm of Zn = 0.012 g / 100 g x 10 6

= 120 ppm